# Optimal Treatment Allocation Strategies for A/B Testing in Partially Observable Time Series Experiments

### v2

@@ 6 Discussion
At the core of our asymptotic theories is the proposed small signal asymptotic framework, which substantially simplifies the asymptotic calculations in time series experiments. More recently, Kuang and Wager, (2023) introduced a weak signal asymptotic in a different context for solving multi-armed bandit problems. The main differences include: (i) Our small signal condition requires the ATE to decay to zero at an arbitrary rate, whereas Kuang and Wager, (2023) requires the difference in mean outcomes between different arms to decay to zero at a more restrictive parametric rate. (ii) Unlike our framework, designed to simplify the asymptotic analysis, their theoretical framework is developed to derive a diffusion convergence limit theorem for sequentially randomized Markov experiments.
@@
Additionally, Farias et al., (2022) and Wen et al., (2024) imposed similar small signal conditions in the same context as ours for A/B testing in time series experiments, aiming either to derive more efficient ATE estimators methodologically or to analyze these estimators theoretically. However, their focus on Markovian environments is more restrictive than ours. They also required the Markov state transition function under the two treatments to be small, a condition less interpretable than our requirement for the ATE to approach zero.
@@ A.2 Proof of Theorem 2 in Controlled VARMA
$=\lim_{T\rightarrow+\infty}\frac{4}{(1-\xi_{\pi}^{2})^{2}T}\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1}{\mbox{Var}}\Big{[}\sum_{t}(U_{t}-\xi_{\pi})\mathbf{Z}_{t}\Big{]}(\mathbb{I}-\mathbf{A})^{-1}e.$
@@ A.3 Proof of Theorem 3: Optimality Conditions for Optimal Design
$\text{Cov}(AC,BD)=\text{Cov}(A,B)\text{Cov}(C,D)+\mathbb{E}(A)\mathbb{E}(B)\text{Cov}(C,D)+\mathbb{E}(C)\mathbb{E}(D)\text{Cov}(A,B)$
for any random variables $A,B,C$ and $D$, we have for any $k\leq q$ and $t>k$ that
$\begin{split}{\mbox{Cov}}((U_{t}-\xi_{\pi})Z_{t},(U_{t-k}-\xi_{\pi})Z_{t-k})\to{\mbox{Cov}}(U_{t}-\xi_{\pi},U_{t-k}-\xi_{\pi})\text{Cov}(Z_{t},Z_{t-k})\\ ={\mbox{Cov}}(U_{t},U_{t-k}){\mbox{Cov}}(Z_{t},Z_{t-k}),\end{split}$ (A.2)
as $t\to\infty$, provided the limit $\xi_{\pi}=\lim_{t}U_{t}$ exists.
@@
This proves the balanced condition for the optimal design.
Meanwhile, under any balanced design $\pi$, we have $\text{Cov}(U_{t},U_{t-k})\text{Cov}(Z_{t},Z_{t-k})=\mathbb{E}(U_{t}U_{t-k})\sum_{j=k}^{q}\theta_{j}\theta_{j-k}$. The asymptotic MSE of the resulting ATE estimator can be simplified as:
@@ C.1 Proof in Controlled ARMA($1,q$)
$-\text{ATE}$ ATE $=\left(\begin{array}[]{cc}\frac{2b}{(1-a)^{2}},&\frac{2}{1-a}\end{array}\right)\left(\begin{array}[]{cc}\xi_{uy^{-1}}&1\\ \xi_{y^{-1}y^{-q-1}}&\xi_{uy^{-q-1}}\\ \end{array}\right)^{-1}\left(\begin{array}[]{c}\displaystyle\frac{1}{T-q}\sum_{t=q+1}^{T}U_{t}Z_{t}\\ \displaystyle\frac{1}{T-q}\sum_{t=q+1}^{T}Y_{t-q-1}Z_{t}\end{array}\right)+o_{p}(T^{-1/2}),$
@@
$-\text{ATE}$ ATE $=\left(\begin{array}[]{cc}\frac{2b}{(1-a)^{2}},&\frac{2}{1-a}\end{array}\right)\left(\begin{array}[]{cc}0&\quad\xi_{y^{-1}y^{-q-1}}^{-1}\\ 1&0\\ \end{array}\right)\left(\begin{array}[]{c}\displaystyle\frac{1}{T-q}\sum_{t=q+1}^{T}U_{t}Z_{t}\\ \displaystyle\frac{1}{T-q}\sum_{t=q+1}^{T}Y_{t-q-1}Z_{t}\end{array}\right)+o_{p}(T^{-1/2})$ (C.6)
@@
where $\xi_{y^{-1}y^{-q-1}}$ is only a function of $a$ and $\theta$, which is independent of $b$ or the magnitude of the ATE. We then look at the second term in (C.6), i.e., $\frac{2b}{(1-a)^{2}}\xi_{y^{-1}y^{-q-1}}^{-1}\frac{1}{T-q}\sum_{t=q+1}^{T}Y_{t-q-1}Z_{t}$. As $Y_{t-q}$ is uncorrelated with $Z_{t}$, $\frac{1}{T-q}\sum_{t}Y_{t-q-1}Z_{t}$ converges to zero by weak law of large number with the convergence rate $T^{-1/2}$. Therefore, we have $\frac{1}{T-q}\sum_{t}Y_{t-q-1}Z_{t}=O_{p}(T^{-1/2})$. Meanwhile, $2b/(1-a)^{2}\propto\text{ATE}$ (especially $b$ is small), which is denoted by $O(\text{ATE})$. Consequently, the second term in (C.6) is $O(\text{ATE})O_{p}(T^{-1/2})$, which tends to 0 within the small signal asymptotics by letting $\text{ATE}\rightarrow 0$ and $T\rightarrow+\infty$. This indicates that:
@@
$-\text{ATE}$ ATE $=\frac{\left(\begin{array}[]{cc}\frac{2b}{(1-a)^{2}},&\frac{2}{1-a}\end{array}\right)}{(\xi_{y^{-1}y^{-q-1}}-\xi_{uy^{-1}}\xi_{uy^{-q-1}})}\left(\begin{array}[]{cc}-\xi_{uy^{-q-1}}&1\\ \xi_{y^{-1}y^{-q-1}}&-\xi_{uy^{-1}}\end{array}\right)\left(\begin{array}[]{c}\displaystyle\frac{1}{T-q}\sum_{t}U_{t}Z_{t}\\ \displaystyle\frac{1}{T-q}\sum_{t}Y_{t-q-1}Z_{t}\end{array}\right)+o_{p}(T^{-\frac{1}{2}})$
@@ Appendix D Estimation, Asymptotic MSEs, and Efficiency Indicators in Controlled VARMA
As AD, UR, and AT are all balanced designs, i.e., $\xi_{\pi}=0$, we consider the proof in the controlled VARMA model without the intercept term $\bm{\mu}$ in this section.
@@ D.1 Controlled VARMA($1,q$)
which is the direct extension of the asymptotic MSE in Controlled ARMA($1,q$) from (C.5). In particular, since $\mathbb{E}\left[U_{t}\right]=0$ and $\mathbb{E}\left[\mathbf{Z}_{t}\right]=\mathbf{0}_{d}$, we have $\text{Cov}(U_{j}\mathbf{Z}_{j},U_{k}\mathbf{Z}_{k})=\text{Cov}(U_{j},U_{k})\text{Cov}(\mathbf{Z}_{j},\mathbf{Z}_{k})$ when $j-k<q$. As $\lim_{\tau\rightarrow+\infty}\text{Cov}(U_{j},U_{k})=1$, we have $\lim_{\tau\rightarrow+\infty}\text{Cov}(U_{j}\mathbf{Z}_{j},U_{k}\mathbf{Z}_{k})=\text{Cov}(\mathbf{Z}_{j},\mathbf{Z}_{k})=\sum_{i=j-k}^{q}\mathbf{M}_{i}\Sigma\mathbf{M}_{i-(j-k)}$.
@@
$=\left[2\mathbf{b}^{\top}(\mathbb{I}-\mathbf{A}^{\top})^{-1}\otimes\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1},2\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1}\right]\frac{1}{T-q}\left(\bm{\xi}_{1,q}\otimes\mathbb{I}_{d}\right)\left(\begin{array}[]{c}\displaystyle\sum_{t}U_{t}\mathbf{Z}_{t}\\ \displaystyle\text{vec}(\sum_{t}\mathbf{Z}_{t}\mathbf{Y}_{t-q-1}^{\top})\end{array}\right)+o_{p}(T^{-1/2})$
@@
$\lim_{\begin{subarray}{c}T\rightarrow+\infty\\ \text{ATE}\to 0\end{subarray}}\text{MSE}(\sqrt{n}\widehat{\textrm{ATE}}(\pi_{\text{AT}}))$ $=4\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1}\left(\sum_{j=0}^{q}\mathbf{M}_{j}\Sigma\mathbf{M}_{j}+2\sum_{k=1}^{q}\sum_{j=k}^{q}(-1)^{k}\mathbf{M}_{j}\Sigma\mathbf{M}_{j-k}\right)(\mathbb{I}-\mathbf{A})^{-1}\bm{e}$

### v4

@@ 6.3 Theory: Small Signal Asymptotic Framework
Meanwhile, similar assumptions have been proposed in the literature on either A/B testing or other fields to simplify theoretical or methodological development. For instance, Kuang and Wager, (2023) introduced a weak signal asymptotic framework in a different context for solving multi-armed bandit problems. The main differences include: (i) Our small signal condition requires the ATE to decay to zero at an arbitrary rate, whereas Kuang and Wager, (2023) requires the difference in mean outcomes between different arms to decay to zero at a more restrictive parametric rate. (ii) Unlike our framework, which is designed to simplify the asymptotic analysis, their theoretical framework is developed to derive a diffusion convergence limit theorem for sequentially randomized Markov experiments.
@@
Additionally, Farias et al., (2022) and Wen et al., (2024) imposed similar small signal conditions in the same context as ours for A/B testing in time dependent experiments, aiming either to derive more efficient ATE estimators methodologically or to analyze these estimators theoretically. However, their focus on Markovian environments is more restrictive than ours. They also required the difference of Markov state transition probabilities under the two treatments to be small, a condition less interpretable than our requirement for the ATE to approach zero. Viviano et al., (2023) assumed a small peer effect condition (Assumption 3) to handle the spatial spillover effect. Such an assumption shares similar spirits to ours, but is designed to develop the optimal design in spatially dependent experiments.
@@ A.2 Proof of Theorem 2 in Controlled VARMA
$\lim_{T\rightarrow+\infty}\text{MSE}(\sqrt{T}\widehat{\text{ATE}}(\pi))=\lim_{T\rightarrow+\infty}\frac{4}{(1-\xi_{\pi}^{2})^{2}T}\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1}{\mbox{Var}}\Big{[}\sum_{t}(U_{t}-\xi_{\pi})\mathbf{Z}_{t}\Big{]}(\mathbb{I}-\mathbf{A})^{-1}e.$
@@ A.3 Proof of Theorem 3: Optimality Conditions for Optimal Design
$\text{Cov}(AC,BD)=\text{Cov}(A,B)\text{Cov}(C,D)+\mathbb{E}(A){E}(B)\text{Cov}(C,D)+\mathbb{E}(C)\mathbb{E}(D)\text{Cov}(A,B)$
for random variables $A,B,C$ and $D$, where $A$is independent of $C$ and $D$, and $B$ is independent of $C$ and $D$, we have for any $k\leq q$ and $t>k$ that
${\mbox{Cov}}((U_{t}-\xi_{\pi})Z_{t},(U_{t-k}-\xi_{\pi})Z_{t-k})&={\mbox{Cov}}(U_{t}-\xi_{\pi},U_{t-k}-\xi_{\pi})\text{Cov}(Z_{t},Z_{t-k})\\ &={\mbox{Cov}}(U_{t},U_{t-k}){\mbox{Cov}}(Z_{t},Z_{t-k}),\end{aligned}$
as $t\to\infty$, provided the limit $\xi_{\pi}=\lim_{t}U_{t}$ exists. The above equation holds as we consider the observation-agnostic design, where $U_{t}$ is independent of $Z_{t}$.
@@
This proves the balanced condition for the optimal design. Meanwhile, under any balanced design $\pi$, we have
$\text{Cov}(U_{t},U_{t-k})\text{Cov}(Z_{t},Z_{t-k})=\mathbb{E}(U_{t}U_{t-k})\sum_{j=k}^{q}\theta_{j}\theta_{j-k}.$
The asymptotic MSE of the resulting ATE estimator can be simplified as:
@@ C.1 Proof in Controlled ARMA($1,q$)
$-\text{ATE}$ ATE $=\left(\begin{array}[]{cc}\frac{2b}{(1-a)^{2}},&\frac{2}{1-a}\end{array}\right)\left(\begin{array}[]{cc}\xi_{uy^{-1}}&1\\ \xi_{y^{-1}y^{-q-1}}&\xi_{uy^{-q-1}}\\ \end{array}\right)^{-1}\left(\begin{array}[]{c}\displaystyle\frac{1}{T-q}\sum_{t=q+1}^{T}U_{t}Z_{t}\\ \displaystyle\frac{1}{T-q}\sum_{t=q+1}^{T}Y_{t-q-1}Z_{t}\end{array}\right)$
$+o_{p}(T^{-1/2}),$
@@
$-\text{ATE}$ ATE (C.6)
$=\left(\begin{array}[]{cc}\frac{2b}{(1-a)^{2}},&\frac{2}{1-a}\end{array}\right)\left(\begin{array}[]{cc}0&\quad\xi_{y^{-1}y^{-q-1}}^{-1}\\ 1&0\\ \end{array}\right)\left(\begin{array}[]{c}\displaystyle\frac{1}{T-q}\sum_{t=q+1}^{T}U_{t}Z_{t}\\ \displaystyle\frac{1}{T-q}\sum_{t=q+1}^{T}Y_{t-q-1}Z_{t}\end{array}\right)+o_{p}(T^{-1/2})$
@@
where $\xi_{y^{-1}y^{-q-1}}$ is only a function of $a$ and $\theta$, which is independent of $b$ or the magnitude of the ATE. We then look at the second term in (LABEL:eq:proof_UR), i.e., $\frac{2b}{(1-a)^{2}}\xi_{y^{-1}y^{-q-1}}^{-1}\frac{1}{T-q}\sum_{t=q+1}^{T}Y_{t-q-1}Z_{t}$. As $Y_{t-q}$ is uncorrelated with $Z_{t}$, $\frac{1}{T-q}\sum_{t}Y_{t-q-1}Z_{t}$ converges to zero by weak law of large number with the convergence rate $T^{-1/2}$. Therefore, we have $\frac{1}{T-q}\sum_{t}Y_{t-q-1}Z_{t}=O_{p}(T^{-1/2})$. Meanwhile, $2b/(1-a)^{2}\propto\text{ATE}$ (especially $b$ is small), which is denoted by $O(\text{ATE})$. Consequently, the second term in (LABEL:eq:proof_UR) is $O(\text{ATE})O_{p}(T^{-1/2})$, which tends to 0 within the small signal asymptotics by letting $\text{ATE}\rightarrow 0$ and $T\rightarrow+\infty$. This indicates that:
@@
$-\text{ATE}$ ATE
$\frac{\left(\begin{array}[]{cc}\frac{2b}{(1-a)^{2}},&\frac{2}{1-a}\end{array}\right)}{(\xi_{y^{-1}y^{-q-1}}-\xi_{uy^{-1}}\xi_{uy^{-q-1}})}\left(\begin{array}[]{cc}-\xi_{uy^{-q-1}}&1\\ \xi_{y^{-1}y^{-q-1}}&-\xi_{uy^{-1}}\end{array}\right)\left(\begin{array}[]{c}\displaystyle\frac{1}{T-q}\sum_{t}U_{t}Z_{t}\\ \displaystyle\frac{1}{T-q}\sum_{t}Y_{t-q-1}Z_{t}\end{array}\right)+o_{p}(T^{-\frac{1}{2}})$
@@ Appendix D Estimation, Asymptotic MSEs, and Efficiency Indicators in Controlled VARMA
As AD, UR, and AT are all balanced designs, i.e., $\xi_{\pi}=0$, we consider the proof in the controlled VARMA model without the intercept term $\bm{\mu}$ in this section. Additionally, we exclude the exogenous variable as well since it remains unaffected by the treatments.
@@ D.1 Controlled VARMA($1,q$)
which is the direct extension of the asymptotic MSE in Controlled ARMA($1,q$) from (C.5). In particular, since $\mathbb{E}\left[U_{t}\right]=0$ and $\mathbb{E}\left[\mathbf{Z}_{t}\right]=\mathbf{0}_{d}$, we have $\text{Cov}(U_{j}\mathbf{Z}_{j},U_{k}\mathbf{Z}_{k})=\text{Cov}(U_{j},U_{k})\text{Cov}(\mathbf{Z}_{j},\mathbf{Z}_{k})$ when $j-k<q$. As $\lim_{\tau\rightarrow+\infty}\text{Cov}(U_{j},U_{k})=1$, we have
$\lim_{\tau\rightarrow+\infty}\text{Cov}(U_{j}\mathbf{Z}_{j},U_{k}\mathbf{Z}_{k})=\text{Cov}(\mathbf{Z}_{j},\mathbf{Z}_{k})=\sum_{i=j-k}^{q}\mathbf{M}_{i}\Sigma\mathbf{M}_{i-(j-k)}.$
@@
$=\left[2\mathbf{b}^{\top}(\mathbb{I}-\mathbf{A}^{\top})^{-1}\otimes\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1},2\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1}\right]\frac{1}{T-q}\left(\bm{\xi}_{1,q}\otimes\mathbb{I}_{d}\right)\left(\begin{array}[]{c}\displaystyle\sum_{t}U_{t}\mathbf{Z}_{t}\\ \displaystyle\text{vec}(\sum_{t}\mathbf{Z}_{t}\mathbf{Y}_{t-q-1}^{\top})\end{array}\right)$
$+o_{p}(T^{-1/2})$
@@
$\lim_{\begin{subarray}{c}T\rightarrow+\infty\\ \text{ATE}\to 0\end{subarray}}\text{MSE}(\sqrt{n}\widehat{\textrm{ATE}}(\pi_{\text{AT}}))$
$=4\bm{e}^{\top}(\mathbb{I}-\mathbf{A})^{-1}\left(\sum_{j=0}^{q}\mathbf{M}_{j}\Sigma\mathbf{M}_{j}+2\sum_{k=1}^{q}\sum_{j=k}^{q}(-1)^{k}\mathbf{M}_{j}\Sigma\mathbf{M}_{j-k}\right)(\mathbb{I}-\mathbf{A})^{-1}\bm{e}$